a dirtbike tire is accelerating a merry-go-round, which starts from the rest, with a constant angular acceleration of a= 2rad/s2 .After 5.3 sec, the person holding onto the merry-go-round, goes flying off. What is the final angular velocity when the person flies off


1
Expert's answer
2021-02-28T07:40:01-0500
"\omega=\omega_0+\alpha t=0+2.0\ \dfrac{rad}{s^2}\cdot 5.3\ s=10.6\ \dfrac{rad}{s}."